OpenStudy (yacoub1993):

what does slashed(/) 0 stand for

OpenStudy (kc_kennylau):

context plz? :)

OpenStudy (yacoub1993):

0 with / on it

OpenStudy (skullpatrol):

the empty set

OpenStudy (kc_kennylau):

oh

OpenStudy (kc_kennylau):

lol @skullpatrol you're quite imaginative \(\phi\) :D

OpenStudy (yacoub1993):

i mean this one ∅

OpenStudy (kc_kennylau):

lol they're just the same but different font

OpenStudy (anonymous):

It's a cancel sign

OpenStudy (kc_kennylau):

https://en.wikipedia.org/wiki/Phi#Use_as_a_symbol It has many use, for example it represents the golden ratio \(\dfrac{1+\sqrt5}2\)

OpenStudy (kc_kennylau):

In D&T it represents diameter?

OpenStudy (yacoub1993):

so for a question like this what would be the answer Use Gaussian elimination to find the complete solution to the system of equations, or state that none exists. 4x - y + 3z = 12 x + 4y + 6z = -32 5x + 3y + 9z = 20 A. {(8, -7, -2)} B. {(-8, -7, 9)} C. ∅ D. {(2, -7, -1)}

OpenStudy (skullpatrol):

who's imaginative now?

OpenStudy (kc_kennylau):

Then that is an empty set, which means that the set of the solutions is empty, i.e. no solution

OpenStudy (skullpatrol):

the set with no members

OpenStudy (yacoub1993):

i got all zeros but the answer choice doenst show all zero

OpenStudy (kc_kennylau):

all zeros ain't the answer, if you plug in all zeros the equations are invalid

OpenStudy (phi):

if you got "all zeros" you got x=0, y=0 , and z=0 but if we use 0 for x,y,z in the equations, for example, the first equation 4x - y + 3z = 12 we get 4*0 - 0 +3*0 which is 0, and 0 ≠ 12 so that can't be the answer

OpenStudy (yacoub1993):

X=0 Y=0 Z=0

OpenStudy (yacoub1993):

@phi X=0 Y=0 Z=0

OpenStudy (phi):

X=0 Y=0 Z=0 is NOT the answer. remember, that if you use these numbers in each equation, you should find the left side of the equation equals the right side. It won't.

OpenStudy (skullpatrol):

the two sides of the equation must have the same value for the equation to be a true statement